# Equal contribution    Corresponding author

### v3

@@ G T2I diffusion model
Appendix 0.B Experimental Environments and Settings
The parameters and devices used for sampling and training are displayed in Tab. 5.
@@ Appendix 0.B Experimental Environments and Settings
SDXL Resolution $1024\times 1024$
@@
SDXL Batch Size 64
@@
SDXL Learning Rate 0.000001
@@
Appendix 0.C Efficiency of AGFSync Illustration
@@
Tab. 6 presents a comprehensive overview of the time required to apply our methodology AGFSync across various stages and models, specifically delineating the durations for tasks such as prompt and QA generation, image generation, VQA scoring, combined CLIP and Aesthetic scoring, and model training. We compare these processes across SD v1.4, SD v1.5, and SDXL models. For prompt and QA generation, all three models require a uniform duration of 12 hours. Image generation and VQA scoring demonstrate variability, with SD v1.4 and SD v1.5 completing in 6 and 7 hours respectively, which contrasts with SDXL’s longer durations of 13 and 12 hours for these tasks. The evaluation of CLIP and Aesthetic scores takes a relatively shorter time, consistently taking 44 minutes across all models. Training times show a distinction between the models, with SD v1.4 and SD v1.5 requiring only 1.5 hours, whereas SDXL necessitates a longer commitment of 3 hours. Tab. 6 underscores the efficiency and resource requirements of our method when applied to different models, providing insightful benchmarks for planning and resource allocation.
Appendix 0.D Complementary Experiments
0.D.1 SDXL Refiner Preformence in TIFA Benchmark
@@
We also applied our method to the SDXL+Refiner model, conducting 40 inference steps on the SDXL model followed by 10 inference steps using the Refiner model. Utilizing the same random seed, the results on the TIFA benchmark are shown in Tab. 7.
@@
Table 7: Results of SDXL+Refiner and AGFSync (Ours)+SDXL+Refiner for VQA score and Aesthetic score on the TIFA benchmark. Red indicates improvement, while Green indicates a decrease. The best scores for each model type are highlighted in Bold. Column "Sum" denotes the sum of improvements on $s_{\text{VQA}}$ and $s_{\text{Aes.}}$
@@
0.D.2 Analysis of Prompts Filtering for High-Quality Image Generation
We also employed the method described in "DreamSync" to filter prompts capable of generating high-quality images from our generated prompts. Specifically, we filtered prompts that could generate images with a VQA score > 0.9 and an Aesthetic score > 0.6 using the SD v1.5 model. Tab. 8 presents the attributes of the text and the questions generated from the filtered prompts, while Tab. 9 displays the number and proportion of each category of prompts obtained through filtering. The results indicate that, although there is no significant change in the nature of the text and the generated questions of the filtered prompts, the proportion of filtered prompts varies greatly across different categories, with the difference in filtering proportion reaching up to 33.5%. This suggests that the types of prompts that the SD v1.5 model, or T2I models in general, excel at vary significantly across categories. Merely selecting prompts capable of producing high-quality images for training is insufficient for a comprehensive approach.
@@ 0.D.2 Analysis of Prompts Filtering for High-Quality Image Generation
0.D.3 Results of Training Models with Different Amounts of Data
Our preference dataset construction method achieved a 100% data conversion efficiency, highlighting the importance of high data conversion efficiency. To further demonstrate this, we randomly selected different proportions of data from our dataset for experiments. Moreover, following the strategy introduced in "DreamSync", we filtered prompts capable of generating high-quality images based on two sets of thresholds (VQA score > 0.85 and Aesthetic score > 0.5, and VQA score > 0.9 and Aesthetic score > 0.6) to construct a preference dataset for training, using SD v1.5 as the base model and conducting the experiment on the TIFA benchmark. The experimental results are shown in LABEL:tab:different_rate_data_train_result. The results indicate that although the Aesthetic scores surpassed the scenario of using all data when trained with 60% and 80% of the data, considering the CLIP score, VQA score, and Aesthetic score together, the larger the volume of data, the better the overall performance of the model. Especially, the improvement in model performance when increasing data usage from 60% to 100% is significantly greater than that from 0% to 60%. Notably, with only our limited constructed preference dataset (i.e. 20% of the preference dataset), we had significant improvements in all aspects, further demonstrating AGFSync’s efficiency.
@@ 0.D.3 Results of Training Models with Different Amounts of Data
Appendix 0.E LLM and VLM Instructions Details
0.E.1 Example Instruction to Generate Image Caption
@@
Please open your mind based on the theme "Natural Landscapes: Includes terrain, bodies of water, weather phenomena, and natural scenes." paintings
@@
0.E.2 Instruction to Generate Question and Answer Pairs with Validation
After obtaining a large number of image captions, we also need to break these captions down into Question-Answer (QA) pairs. For this step, we use Gemini Pro, requesting it to decompose each image caption into 15 QA pairs, with each caption processed six times. Finally, we filter out the QA pairs that were generated repeatedly. Below is the instruction given to Gemini Pro for breaking down image captions into QA pairs.
@@ 0.E.2 Instruction to Generate Question and Answer Pairs with Validation
    You also need to indicate whether the question you provided is an invalid question of the "not mentioned in the text" type, with 0 representing an invalid question and 1 representing a minor question.
@@
"question_id": "The number of the issue you generated, starting with 1",
"question": "A general interrogative sentence you derive from breaking down the text should inquire whether the image conforms to the content of the text. The answer to this question must be found based on the text, not on common sense, etc. The answer must not be unmentioned in the text, and according to the text, the answer to this question must be yes’.",
"answer": "The real answer to the question according to the text provided. The answer should be yes’",
"element_type": "The type of problem. (object, human, animal, food, activity, attribute, counting, color, material, spatial, location, shape, other)",
"element": "The elements mentioned in the question, or the specific elements asked by the question",
"flag": "Check if the correct answer to the question you generated is an invalid question such as not mentioned, with 0 being an invalid question and 1 being not an invalid question"
# There should be more questions here, because a text should be broken down into multiple questions, and the number of questions is up to you
@@
0.E.3 Detailed Instruction When Using GPT-4V for Evaluation
@@
"reason": "your reason",
"choice": "A/B"
@@
"reason": "your reason",
"choice": "A/B"
@@
"reason": "your reason",
"choice": "A/B"
Appendix 0.F Details of Generated Prompts and Preference Dataset
0.F.1 Statistics of the AI-Generated Captions
Tab. 11 presents more detailed data of our generated dataset. In addition to this, in our generated QA pairs, the counts for each category are as follows: shape (2385), counting (3809), material (4495), food (4660), animal (5533), color (12749), human (17921), spatial (21878), other (24513), location (42914), object (77783), activity (83712), and attribute (101713).
@@ 0.F.1 Statistics of the AI-Generated Captions
0.F.2 Example Generated Image Caption and Corresponding QA Pairs
For the example prompt: A vast, open savannah, where golden grasses sway in the wind, dotted with acacia trees and herds of majestic elephants and giraffes, as the sun sets on the horizon. We have the corresponding QA pairs as in Tab. 12. As indicated in Tab. 12, for each generated question, we require the LLM to provide the main element and type of element being asked in the question.
@@ 0.F.2 Example Generated Image Caption and Corresponding QA Pairs
0.F.3 Example Preference Dataset Generated by AGFSync

### v6

@@ G T2I diffusion model
Appendix B Limitations
Firstly, AGFSync relies on existing large language models (LLMs) and aesthetic scoring models, whose performance and accuracy could be influenced by the biases and limitations of the LLMs. Secondly, while we introduce random noise to increase image diversity, this method might lead to a reduction in consistency between some images and their text prompts. In addition, due to the high cost of time or money, we have not adopted LLaVA (Liu et al. 2023), GPT-4V and latest advanced multimodal large models (Qin et al. 2023; Liu et al. 2023; Dai et al. 2023a; Zhu et al. 2023; Ye et al. 2023; Zhou et al. 2024b) to generate prompts or QA pairs. In terms of image evaluation, we employ VQA scores, CLIP scores, and aesthetic scores, which may not capture all aspects of image quality.
Appendix C More Visualized Results
Fig. 5 showcases the effectiveness of AGFSync by comparing text-to-image generation results before and after applying the algorithm to SDXL. Through these side-by-side comparisons, we can observe improvements in both prompt faithfulness and image aesthetics, demonstrating how AGFSync enhances the model’s capabilities without requiring human intervention.
Figure 5: We introduce AGFSync: a model-agnostic training algorithm that improves text-to-image (T2I) generation models’ faithfulness and coherence to text inputs and image aesthetics without human interventions. The images showcase a comparison of the results before and after finetuning SDXL with AGFSync.
Appendix D Experimental Environments and Settings
The softwares, more detailed hyperparameters and devices used for sampling and training are displayed in Table 6.
@@ Appendix D Experimental Environments and Settings
Config Detail
Operating System Ubuntu 20.04
Python version 3.10
PyTorch version 2.2.0
transformers version 4.37.2
diffusers version 0.25.0
@@
SDXL-base Resolution $1024\times 1024$
@@
SDXL-base Batch Size 64
@@
SDXL-base Learning Rate 0.000001
@@
CPUs Intel(R) Xeon(R) Gold 6248R CPU @ 3.00GHz
Appendix E Efficiency of AGFSync Illustration
@@
Table 7 presents a comprehensive overview of the time required to apply our methodology AGFSync across various stages and models, specifically delineating the durations for tasks such as prompt and Q&A generation, image generation, VQA scoring, combined the CLIP score and aesthetic score, and model training. We compare these processes across SD v1.4, SD v1.5, and SDXL-base models. For prompt and Q&A generation, all three models require a uniform duration of 12 hours. Image generation and VQA scoring demonstrate variability, with SD v1.4 and SD v1.5 completing in 6 and 7 hours respectively, which contrasts with SDXL-base’s longer durations of 13 and 12 hours for these tasks. The evaluation of CLIP and aesthetic scores takes a relatively shorter time, consistently taking 44 minutes across all models. Training times show a distinction between the models, with SD v1.4 and SD v1.5 requiring only 1.5 hours, whereas SDXL-base necessitates a longer commitment of 3 hours. Table 7 underscores the efficiency and resource requirements of our method when applied to different models, providing insightful benchmarks for planning and resource allocation.
Appendix F Discussion on the Marginal Improvement of CLIP Score
In the ablation study (see in Table 4), we observed that the improvement in CLIP score was relatively small. To better understand this, we conducted a thorough analysis and provide a detailed explanation of the marginal improvement in CLIP score.
Firstly, each pair of images in the preference dataset is generated by the same model, which leads to relatively small differences between the two images in each data point. CLIP score is calculated by encoding both the image and the text as embeddings and measuring their similarity, reflecting the overall alignment between the image and the text. This method results in smaller differences in the CLIP score between the two images in the preference dataset. As mentioned in the main text, we scale CLIP score, VQA score, and aesthetic score to a 0-100 range. Statistical analysis of the preference dataset shows that the difference in CLIP score between “good” and “bad” images is, on average, 5.34, while the differences in VQA score and aesthetic score are 27.49 and 19.48, respectively. Therefore, after training the model, the improvement in CLIP score is somewhat smaller compared to VQA score and aesthetic score.
To further investigate this, we conducted ablation experiments, the results of which are shown in Table 4. When using CLIP score alone for evaluation, the improvement is 0.2; when using all three scores (CLIP score, VQA score, and aesthetic score) together, the improvement in CLIP score is 0.3. Notably, when CLIP score is excluded from the evaluation, the improvement is less than 0.1.
Despite the relatively small improvement in CLIP score, it still plays an important role. CLIP score provides a semantic-level evaluation of images and its continuous output compensates for the limitations of VQA score, which is a discrete value. Therefore, even though the improvement in CLIP score is relatively modest, we still consider it an essential part of the evaluation. As shown in our ablation experiments, the inclusion of CLIP score leads to significant improvements in both VQA score and aesthetic score, further validating the importance of CLIP score in multidimensional evaluation.
In conclusion, while the improvement in CLIP score is marginal, it provides additional semantic alignment information in the evaluation process. Especially when combined with VQA and aesthetic scores, it significantly enhances the overall evaluation. Therefore, we believe that CLIP score remains a valuable component of our evaluation methodology.
Appendix G Discussion on the Weight Selection Method
In this study, we optimized the weights of CLIP score, VQA score, and aesthetic score through grid search to obtain a weighted score. The selection of these weights is crucial as it directly influences the overall performance of the model and the contribution of each scoring dimension to the final result. We employed the following approach for weight selection:
G.1 Grid Search Optimization
To ensure that the total sum of the weights for each score equals 1 and to avoid excessively large or small weights for any score, we set upper and lower bounds for each score’s weight and selected specific candidate values within these ranges. This strategy effectively reduces the computational cost of the grid search. Specifically, the candidate values for the weight of CLIP score are $\{0.3,0.35,0.4,0.45,0.5,0.55,0.6\}$, for VQA score are $\{0.3,0.35,0.4,0.45,0.5,0.55,0.6\}$, and for aesthetic score are $\{0.05,0.1,0.15,0.2\}$. We ensured that the sum of the weights for the three scores always equaled 1. Using this approach, we are able to find the optimal combination among multiple possible weight configurations.
G.2 Reason for the Small Weight of aesthetic Score
Since the primary goal of this study is to improve the consistency between the generated images and the input text in the image generation model, we found that the weight of the aesthetic score should be relatively small. In the experiments, we also discovered that when the weight of aesthetic score is too large, it may slightly damage the consistency between the image and text. Therefore, during the grid search process, we set the weight range for aesthetic score to be smaller to ensure greater improvement in text-image consistency.
G.3 Setting Weights for Other Scores
Both CLIP score and VQA score are closely related to text-image consistency. To emphasize consistency, we set the weights of these two scores higher. Through grid search, we identified a weight configuration for CLIP score, VQA score, and aesthetic score that significantly improved consistency while also enhancing the overall generation quality.
G.4 Ablation Study Validation
During the experiments, we conducted an ablation study (see Table 4) to validate the effectiveness of different weight configurations. The results show that training with all three scores significantly outperforms training with any single score or a combination of two scores. This further demonstrates the effectiveness of the chosen weight configuration in improving both consistency and generation quality.
In summary, the weight selection is based on the model’s goal—improving the consistency between text and image. As such, higher weights are assigned to CLIP score and VQA score, both of which are related to consistency, while aesthetic score was given a relatively lower weight. Through grid search optimization, we ultimately selected the most effective weight combination, which resulted in significant improvements in the generation quality.
Appendix H Complementary Experiments
H.1 SDXL Refiner Preformence in TIFA Benchmark
@@
We also applied our method to the SDXL+Refiner model, conducting 40 inference steps on the SDXL-base model followed by 10 inference steps using the Refiner model. Utilizing the same random seed, the results on the TIFA benchmark are shown in Table 8.
@@
Table 8: Results of SDXL+Refiner and AGFSync (Ours)+SDXL+Refiner for VQA score and aesthetic score on the TIFA benchmark. Red indicates improvement, while Green indicates a decrease. The best scores for each model type are highlighted in Bold. Column “Sum” denotes the sum of improvements on $s_{\text{VQA}}$ and $s_{\text{Aes.}}$
@@
H.2 Comparison of Win Rates with Draw Thresholds on HPS v2 Benchmark
Table 9: Win & draw rates under gap thresholds of applying AGFSync on base model vs the original base model on HPS v2.
Tie Threshold Score SD v1.4 SD v1.5 SDXL-base
Win Draw Win Draw Win Draw
0.1 CLIP 52.5% 6.9% 55.4% 4.8% 55.6% 4.3%
Aes. 56.8% 4.9% 63.3% 2.9% 76.3% 1.9%
0.01 CLIP 56.0% 0.6% 57.8% 0.4% 57.6% 0.3%
Aes. 59.1% 0.6% 64.3% 0.3% 77.1% 0.3%
0.001 CLIP 56.2% 0.1% 57.9% 0.03% 57.7% 0.0%
Aes. 59.2% 0.03% 64.6% 0.0% 77.3% 0.05%
In the HPS v2 benchmark test, we establish a gap threshold (0.1, 0.01, 0.001) to determine the outcomes of comparisons. Results are considered a draw when the absolute value of the gap is less than or equal to this threshold. In Table 9, we present the win rates and probabilities of drawing at different gap thresholds for the SD v1.4, SD v1.5, and SDXL-base models trained using AGFSync, compared with the original models, across both CLIP scores and aesthetic scores. The results clearly demonstrate that the models finetuned with AGFSync consistently outperform the base models. Moreover, even when adjusting the gap threshold to 0.1, the changes in win rates remain minimal.
H.3 Analysis of Prompts Filtering for High-Quality Image Generation
We also employ the method described in “DreamSync” to filter prompts capable of generating high-quality images from our generated prompts. Specifically, we filter prompts that can generate images with a VQA score >0.9 and an aesthetic score >0.6 using the SD v1.5 model. Table 10 presents the attributes of the text and the questions generated from the filtered prompts, while Table 11 displays the number and proportion of each category of prompts obtained through filtering. The results indicate that, although there is no significant change in the nature of the text and the generated questions of the filtered prompts, the proportion of filtered prompts varies greatly across different categories, with the difference in filtering proportion reaching up to 33.5%. This suggests that the types of prompts that the SD v1.5 model, or T2I models in general, excel at vary significantly across categories. Merely selecting prompts capable of producing high-quality images for training is insufficient for a comprehensive approach.
@@ H.3 Analysis of Prompts Filtering for High-Quality Image Generation
H.4 Results of Training Models with Different Amounts of Data
Our preference dataset construction method achieves a 100% data conversion efficiency, highlighting the importance of high data conversion efficiency. To further demonstrate this, we randomly select different proportions of data from our dataset for experiments. Moreover, following the strategy introduced in ”DreamSync”, we filter prompts capable of generating high-quality images based on two sets of thresholds (VQA score >0.85 and aesthetic score >0.5, and VQA score >0.9 and aesthetic score >0.6) to construct a preference dataset for training, using SD v1.5 as the base model and conducting the experiment on the TIFA benchmark. The experimental results are shown in LABEL:tab:different_rate_data_train_result. The results indicate that although the aesthetic scores surpass the scenario of using all data when trained with 60% and 80% of the data, considering the CLIP score, VQA score, and aesthetic score together, the larger the volume of data, the better the overall performance of the model. Especially, the improvement in model performance when increasing data usage from 60% to 100% is significantly greater than that from 0% to 60%. Notably, with only our limited constructed preference dataset (i.e. 20% of the preference dataset), we have significant improvements in all aspects, further demonstrating AGFSync’s efficiency.
@@ H.4 Results of Training Models with Different Amounts of Data
H.5 Prompt Utilization Rate Analysis
Figure 6: In our generated dataset constructed with images generated by SDXL-base, the proportion of prompts that can be filtered out based on varying thresholds.
To compare with DreamSync (Sun et al. 2023), in Fig. 6, we present the number of prompts that can be selected from our generated dataset as the thresholds for VQA score and aesthetic score vary. Specifically, these are prompts for which at least one generated image scores above the respective thresholds. With DreamSync’s 0.9 and 0.6 thresholds for two metrics, only 48.8% of prompts satisfying both, meaning a low data conversion efficiency of 48.8% on our dataset. However, AGFSync’s approach merely requires selecting the best and worst images without imposing any threshold constraints, thereby achieving a data conversion efficiency of 100%.
Appendix I LLM and VLM Instructions Details
I.1 Example Instruction to Generate Image Caption
@@
Please open your mind based on the theme ”Natural Landscapes: Includes terrain, bodies of water, weather phenomena, and natural scenes.” paintings
@@
I.2 Instruction to Generate Question and Answer Pairs with Validation
After obtaining a large number of image captions, we also need to break these captions down into Question-Answer (QA) pairs. For this step, we use Gemini Pro, requesting it to decompose each image caption into 15 QA pairs, with each caption processed six times. Finally, we filter out the QA pairs that are generated repeatedly. Below is the instruction given to Gemini Pro for breaking down image captions into QA pairs.
@@ I.2 Instruction to Generate Question and Answer Pairs with Validation
    You also need to indicate whether the question you provided is an invalid question of the ”not mentioned in the text” type, with 0 representing an invalid question and 1 representing a minor question.
@@
1[
2 {
3 "question_id": "The number of the issue you generated, starting with 1",
4 "question": "A general interrogative sentence you derive from breaking down the text should inquire whether the image conforms to the content of the text. The answer to this question must be found based on the text, not on common sense, etc. The answer must not be unmentioned in the text, and according to the text, the answer to this question must be yes’.",
5 "answer": "The real answer to the question according to the text provided. The answer should be yes’",
6 "element_type": "The type of problem. (object, human, animal, food, activity, attribute, counting, color, material, spatial, location, shape, other)",
7 "element": "The elements mentioned in the question, or the specific elements asked by the question",
8 "flag": "Check if the correct answer to the question you generated is an invalid question such as not mentioned, with 0 being an invalid question and 1 being not an invalid question"
9 }
10 # There should be more questions here, because a text should be broken down into multiple questions, and the number of questions is up to you
11]
@@
I.3 Detailed Instruction When Using GPT-4V for Evaluation
@@
1{
2 "reason": "your reason",
3 "choice": "A/B"
4}
@@
1{
2 "reason": "your reason",
3 "choice": "A/B"
4}
@@
1{
2 "reason": "your reason",
3 "choice": "A/B"
4}
Appendix J Details of Generated Prompts and Preference Dataset
J.1 Statistics of the Prompts for Different Categories
The statistics of our LLM-generated prompts for preference candidate sets’ image captions, comprising 45,834 prompts, are presented in Table 13 across 12 distinct categories: Natural Landscapes, Cities and Architecture, People, Animals, Plants, Food and Beverages, Sports and Fitness, Art and Culture, Technology and Industry, Everyday Objects, Transportation, and Abstract and Conceptual Art.
Table 13: Distribution of prompts across different categories in our dataset.
Category Count
Natural Landscapes 5,733
Cities and Architecture 6,291
People 5,035
Animals 3,089
Plants 4,276
Food and Beverages 3,010
Sports and Fitness 2,994
Art and Culture 2,432
Technology and Industry 3,224
Everyday Objects 2,725
Transportation 4,450
Abstract and Conceptual Art 2,575
Total 45,834
As shown in Table 13, the distribution of prompts across categories varies, with Cities and Architecture having the highest count (6,291) and Art and Culture having the lowest (2,432). This diversity in prompt distribution ensures a wide range of concepts and subjects for our T2I model to learn from and generate images.
J.2 Statistics of the AI-Generated Captions
Table 14 presents more detailed data of our generated dataset. In addition to this, in our generated QA pairs, the counts for each category are as follows: shape (2385), counting (3809), material (4495), food (4660), animal (5533), color (12749), human (17921), spatial (21878), other (24513), location (42914), object (77783), activity (83712), and attribute (101713).
@@
J.3 Example Generated Image Caption and Corresponding QA Pairs
For the example prompt: A vast, open savannah, where golden grasses sway in the wind, dotted with acacia trees and herds of majestic elephants and giraffes, as the sun sets on the horizon. We have the corresponding QA pairs as in Table 15. As indicated in Table 15, for each generated question, we require the LLM to provide the main element and type of element being asked in the question.
@@
J.4 Example Preference Dataset Generated by AGFSync